OpenStudy (anonymous):

integral (6(sinx)(cos^2(x)) dx

OpenStudy (anonymous):

tailor made for u - sub. put u = cos(x), du = -sin(x)dx and get \[-6\int u^2 du\]

OpenStudy (anonymous):

anti derivative is \[-\frac{6}{3}u^3=-2cos^3(x)\]

OpenStudy (anonymous):

+C if you need it

OpenStudy (anonymous):

thanks!